OpenStudy (anonymous):

Martha can rake the leaves in her yard in 2hr. Her brother can do the job in 4hr. How long will it take them to do it together?

OpenStudy (anonymous):

\[\frac{2\times 4}{2+4}\] \[\frac{8}{6}=\frac{4}{3}\]

OpenStudy (anonymous):

one and one third hour, aka one hour twenty minutes

OpenStudy (anonymous):

Can you go look at the anwers on the last question and tell me which one please

OpenStudy (anonymous):

k